### v4

# Multipartite Entanglement Distribution in Quantum Networks using Subgraph Complementations

Abstract
Quantum networks are important for quantum communication and allow for several tasks such as quantum teleportation, quantum key distribution, quantum sensing, and quantum error correction. Graph states are a specific class of multipartite entangled states that can be represented by graphs. We propose a novel approach for distributing graph states across a quantum network. We show that the distribution of graph states can be characterized by a system of subgraph complementations, which we also relate to the minimum rank of the underlying graph and the degree of entanglement quantified by the Schmidt-rank of the quantum state. We analyze resource usage for our algorithm and show that it improves on the number of qubits, bits for classical communication, and EPR pairs utilized, as compared to prior work. In fact, the number of local operations and resource consumption for our approach scales linearly in the number of vertices. This produces a quadratic improvement in completion time for several classes of graph states represented by dense graphs, which translates into an exponential improvement by allowing parallelization of gate operations. This leads to improved fidelities in the presence of noisy operations, as we show through simulation in the presence of noisy operations. Common classes of graph states are classified along with their optimal distribution time using subgraph complementations. We find a close to optimal sequence of subgraph complementation operations to distribute an arbitrary graph state, and establish upper bounds on distribution time along with providing approximate greedy algorithms.
I Introduction
footnotetext: Main contactfootnotetext: Manning College of Information and Computer Sciences
I-A Motivation
The theory of multipartite entanglement has been extensively used to study the distribution of graph states in a quantum network [1, 2, 3]. Creating entangled graph states efficiently is important for the development of quantum networks, having applications in areas such as quantum metrology [4], quantum error correction [5] and distributed quantum computing [6]. Graph states also form the basis of the one-way model of computation for quantum computers, which is universal and operates by performing single qubit measurements on an initially created resource graph state [7].
In this paper, we first consider a quantum network in ideal conditions with no noisy communication and a limited number of perfect quantum memories. We address the problem of distributing a graph state under such conditions and present an algorithm to do so for a network having an arbitrary topology. We are in particular interested in gaining a better understanding of the limitations in resource consumption (of resources such as EPR pairs and number of qubits) and runtime associated with distributing graph states, and how they can be optimized. Secondly, we show that our algorithm leads to improved fidelities of the graph state in the presence of noisy operations.
I-B Related Work
The distribution of arbitrary graph states in a quantum network has received attention recently. Some of the key parameters for distributing graph states are the number of Bell pairs consumed and the number of time steps required. Two approaches similar to our work (i.e. a noiseless network with an arbitrary topology) are provided by [2, 3]. The authors of [2] do not optimize resource consumption for an arbitrary graph state, but provide a framework for future work on this problem by analyzing resource consumption in a noiseless and lossless network. Their algorithm utilizes $(n-1)$ Bell pairs in one time step to distribute an $n$-qubit GHZ state, and $\frac{n(n-1)}{2}$ Bell pairs in $(n-1)$ time steps to distribute an arbitrary graph state with $n$ vertices [2]. Another approach using an identical setting [3] utilizes $\frac{3n^{2}-2n}{8}$ Bell pairs and takes $(n-1)$ time steps to distribute an arbitrary graph state. Both of these approaches assume that multiple gates can be applied on a single qubit in one time step. However, unlike these approaches, we consider the gate-based model of computation in which a qubit can only have one gate applied to it in one time step. We also assume that Bell pairs are distributed in advance between some central node and end nodes. This increases the distribution times of [2] and [3] by a factor of $n$, so they both require $O(n^{2})$ time steps.
Other work [8, 9] considers additional factors such as the probabilistic nature of generating and the time for distributing Bell pairs, accounting for memory decoherence and noisy Bell state measurements. These factors significantly impact the rate and fidelity of distributing a graph state and are of interest to practically implement such protocols in a physical network.
The algorithm provided by Avis et al. [8] (and Fischer, Towsley [3], under the assumption that the physical network has a star topology) improves on the work of [2] by proposing the following approach: 1) Prepare the desired $n$-qubit graph state locally at a central node; 2) distribute Bell pairs between the central node and each of the $n$ end nodes; 3) teleport each qubit in the central node to each of the end nodes, consuming all Bell pairs in the process.
This approach utilizes $n$ Bell pairs, $O(|E|)=O(n^{2})$ time steps, and requires $2n$ qubits to be stored in the central node. We improve upon each one of these factors as described in the following section. Note that the distribution times for some of the above approaches can be improved to $O(n)$ by parallelizing gate operations. We also analyze the effect of incorporating parallelization in our approach and show a comparable speedup (within a $O(\log n)$ factor) for distributing all graph states, and an exponential improvement in distribution time, i.e. $O(\log n)$ time steps, for certain relevant classes of graphs.
I-C Summary of Contributions
We propose a novel graph state distribution algorithm both with and without parallelization and analyze it first for the noiseless case. We then examine the effects of noise through simulations on various graph states. Our protocol is based on a specific operation on graph states, termed a subgraph complementation ($SC$), which should not be confused with local complementation. We then show that any graph state can be distributed in the network through a sequence of $SC$ operations. Along with introducing this protocol, we make the following contributions, improving upon prior work.
  • We show that our graph state distribution algorithm utilizes $(n-1)$ Bell pairs and requires at most $(n-1)$ qubits at the central node. The number of Bell pairs is clearly optimal and the number of qubits at the central node approximately scales with the average degree of the vertices in the graph.
  • We show that our algorithm requires $O(n)$ time steps and $O(n)$ bits of classical communication to distribute sparse graphs, as well as several relevant classes of dense graphs such as complete graphs (GHZ states) and complete bipartite graphs, but may take up to $O(n^{2})$ resources for some dense graphs.
  • By parallelizing gate operations and using auxiliary qubits, we show that the distribution time improves to $O(\log n)$ time steps for the same classes of dense graphs. We obtain a general upper bound of $O(n\log n)$ time steps for arbitrary dense graphs and $O(n)$ for sparse graphs, as well as $O(n)$ time steps for random graphs (sparse and dense) with high probability.
  • We establish a relation between the number of $SC$ operations needed to distribute a graph state |$\ket{G}$ and the entanglement of |$\ket{G}$ (quantified by its average Schmidt-rank over all possible bipartitions).
  • We also carry out simulations (with and without parallelization) of our algorithm in the presence of gate operations characterized by depolarizing noise, and show that our algorithm almost always increases fidelity compared to other approaches.
II Background
Graph states are a class of multipartite entangled states that can be represented by graphs, with each vertex representing a qubit and each edge representing a $CZ$ (controlled-$Z$) gate. A graph state $\left|G\right\rangle$ corresponding to the undirected graph $G=(V,E)$ is defined as
$\left|G\right\rangle=\prod_{(a,b)\in E}CZ_{ab}\left|+\right\rangle^{|V|}\,,$
[…]
We utilize the following graph operations corresponding to quantum gates or measurements to modify a graph state in this paper — edge addition/deletion ($CZ$ gate), local complementation of a vertex ($LC$) and measurement of a qubit in either the $Y$ or $Z$ basis. An $LC$ operation on a vertex complements the induced subgraph formed by the neighborhood of a vertex, whereas a $Z$-basis measurement removes a vertex from the graph [1]. A $Y$-basis measurement is equivalent, up to single qubit Clifford gates, to an $LC$ operation followed by removal of the vertex [1].
We also need an operation denoted as subgraph complementation [10] (or $SC$) on some subset of vertices $V^{\prime}\subseteq V$ of the graph. Subgraph complementation of $V^{\prime}$ corresponds to complementing the induced subgraph on vertices $V^{\prime}$ of the graph $G$. That is, for each pair of vertices $v_{i},v_{j}\in V^{\prime}$, we either add or remove the corresponding edge $(v_{i},v_{j})$.
An $SC$ operation is similar to an $LC$ operation since an $LC$ on a particular vertex $v$ is equivalent to an $SC$ carried out on the neighborhood of $v$. However, $SC$ operations do not correspond to $LC$ operations since they can be on any arbitrary induced subgraph of $G$. The one exception is when the subgraph is the neighborhood of $v$ for some $v\in V$ not in the subgraph. Also, note that $SC$ operations commute.
III Subgraph complementation protocol
We describe the protocol for the setting of no noise and deterministic Bell pair generation. This section will be followed by a discussion of $SC$ operations as part of our algorithm to distribute an arbitrary graph state.
III-A Single subgraph complementation
We first present a protocol for performing a single $SC$ operation on any subset of nodes $V_{\alpha}\subseteq V$ of the network, requiring $O(|V_{\alpha}|)$ time steps. This will serve as a building block for our graph state distribution algorithm.
The protocol by Avis et al. [8] distributes a state by locally creating it and then teleporting it, which is costly in memory usage. Roughly speaking, we get around this by constructing the desired graph state step by step through $SC$ operations which consist of local CZ gates (acting only on $n$ qubits as opposed to the $2n$ qubits in Avis et al. [8]) and local complementations.
More formally, consider a network with a central node $A$ that contains $k-1$ qubits $\{a_{i}:0\leq i<k-1\}$. The network also contains a set of $k-1$ nodes containing one qubit each $\{c_{i}:0\leq i<k-1\}$. Our setup assumes that $(k-1)$ Bell pairs of the form $(a_{i},c_{i}),0\leq i<(k-1)$, are already distributed between the nodes. Note that this defines two network graphs corresponding to the nodes (the physical graph) and qubits (the graph state) respectively.
The following protocol describes how to distribute a $k-clique$, or a fully connected graph of $k$ vertices, among the qubits in the set $\{a_{i}\}\bigcup\{c_{i}\}$ where qubits $a_{i}$ are in the central node $A$ and qubits $c_{i}$ are in the other end nodes. An example operation for $k=4$ can be found in Fig. 1.
Figure 1: 4-clique addition operation. If a 4-clique is the final state to be distributed, qubits $a_{1}$ and $a_{2}$ need to be measured out in the $Z$-basis at the end. Else, preserve the qubits for further $SC$ operations.
A
Subgraph complementation protocol on $V_{\alpha}$
  1. 1.
    $\forall a_{i}:i\neq 0\implies CZ(a_{i},a_{0})$ — Connect all qubits in $A$ with an edge to $a_{0}$. The resulting state corresponds to state (ii) in the figure for 4 qubits.
  2. 2.
    $\forall a_{i}:i\neq 0\implies LC(a_{i})$ — Locally complement each qubit in $A$ except $a_{0}$.
  3. 3.
    $\forall a_{i}:i\neq 0\implies CZ(a_{i},a_{0})$ — This is a repeat of the first step, except this time the edge between $a_{i}$ and $a_{0}$ is deleted. The resulting state corresponds to state (iii).
  4. 4.
    $LC(a_{0})$ — The resulting state corresponds to state (iv). This forms the complete graph for $V_{\alpha}$.
If a $k$-clique (corresponding to a $k$-qubit GHZ state) is the final state we want to distribute, then some qubits may need to be measured out in the $Z$-basis following the protocol (e.g. $a_{1}$ and $a_{2}$ in Fig. 1). However, if we want to distribute an arbitrary graph state, then it is beneficial to actually preserve connections of the form $(a_{i},c_{i})$ between qubits for carrying out additional $SC$ operations, in order to avoid sharing a new set of Bell pairs each time. This is described in more detail below.
III-B Sequence of subgraph complementations
In order to carry out another $SC$ operation after obtaining a state similar to state (iv) in Fig. 1, we need to restore direct connections $(a_{i},c_{i})$ between the central node and each of the end nodes, similar to the connections in state (i) in Fig. 1. We refer to this as an edge reset operation. This simply involves carrying out steps 1–3 in the subgraph complementation protocol in reverse order. This is illustrated in Fig. 2. The only difference between the new state (v) and state (i) is that the end nodes form a clique. The $SC$ protocol can then be carried out on a new set of vertices to add another clique, but since $CZ^{2}=I$, the edges of the clique get added modulo 2 (every two additions of the same edge removes it). More specifically, applying the sequence of operations 1–4 on any set of vertices $V_{\alpha}\subseteq V$ corresponds to performing another $SC$ operation on $V_{\alpha}$. This can be used to successively perform several $SC$ operations.
Finally, after performing some sequence of $SC$ operations, we would like to select which connections (edges) to keep or not to keep between the central node $A$ and the other nodes. Since $A$ is also part of the network, simply perform step 1 followed by step 2 of the $SC$ protocol to add any edges between $A$ (resp. $a_{i}$) and $C_{i}$ ($c_{i}$), and measure out all the other qubits.
Figure 2: Edge reset operation.
III-C Analysis
In order to complement a subgraph of $k$ vertices in a network, this protocol utilizes $k-1$ Bell pairs and $k-1$ qubits in the central node. A total of $(2k-2)$ $CZ$ operations and $k$ $LC$ operations are required by the basic protocol. This will be followed by a linear number of $Z$-measurements (maximum $k$), $LCs$ and $CZs$ (maximum $k-1$ in conjunction with each other). Thus, the number of time steps required to carry out this protocol is $O(k)$.
IV Graph State Distribution Algorithm
In this section, we describe an algorithm for distributing an arbitrary graph state, which uses $SC$ operations as a primitive. Specifically, we show that any graph state can be generated through a sequence of $SC$ operations, and also analyze the distribution time.
IV-A Subgraph complementation system
Any arbitrary graph $G=(V,E)$ with $n$ vertices can be represented by a subgraph complementation system $\mathcal{SC}_{sys}=\{V_{1},\ldots,V_{d}\}$ [10] where $V_{i}$ denotes a single $SC$ operation on $V_{i}\subseteq V$ and $d=|\mathcal{SC}_{sys}|$. Note that $SC$ operations can be carried out in any order, and that this decomposition is not unique. Let the subgraph complementation number $\mathbb{c}_{2}(G)$ denote the minimum cardinality $d$ (over all possible $\mathcal{SC}_{sys}$ for a graph $G$) of such a set [10]. For instance, consider the case of building the wheel graph $W_{5}$ as shown leftmost in Fig. 3. For one possible $\mathcal{SC}_{sys}$, all 8 edges (which are 2-cliques) are added one by one, corresponding to one $SC$ operation each, so $d=8$.
Figure 3: Subgraph complementation system for $W_{5}$ [10].
In the example shown in the figure, only three $SC$ operations (solid lines) are required and $\mathcal{SC}_{sys}=\{\{1,2,4,5\},\{2,3,4\},\{1,3,5\}\}$, where each subset of $\mathcal{SC}_{sys}$ contains the vertices in an $SC$ operation according to the numbering in Fig. 3, thus $d=3$. In fact, we can show that this is the minimum number of $SC$ operations required to build the given graph $W_{5}$. Thus, $\mathbb{c}_{2}(W_{5})=3$.
IV-B Time for subgraph complementation
For our analysis, we assume that Bell pairs are distributed in advance. We consider first the gate-based model of computation with perfectly noiseless quantum channels for communication (see Section VII for a numerical analysis in the presence of noise). Let $T(G,\mathcal{SC}_{sys})$ denote the time required by our protocol to distribute a graph state $\left|G\right\rangle$ according to subgraph complementation system $\mathcal{SC}_{sys}$. Denote by $t(k)$ the time required for a single $SC$ operation on $k$ vertices, $k\in\mathbb{N}$. It follows from the prior analysis section that $t(k)=\gamma k=O(k),$ for some $\gamma\in\mathbb{R^{+}}$, since time is proportional to the number of $CZ$ gates performed. This directly leads to an expression for $T(G,\mathcal{SC}_{sys})$ as stated below.
Theorem IV.1.
Given a graph $G$ with $n$ vertices and the corresponding system $\mathcal{SC}_{sys}$ with minimal size $\mathbb{c}_{2}(G)$ such that each $SC$ operation on $V_{i}$ is performed on $k_{i}$ vertices, then $T(G,\mathcal{SC}_{sys})=\mathbb{c}_{2}(G)\cdot t(\overline{k})\,,$ where $\overline{k}=\frac{1}{d}\sum_{i=1}^{d}k_{i}$.
Proof.
Given a graph $G$ and an $SC$ system $\mathcal{SC}_{sys}$,
$T(G,\mathcal{SC}_{sys})=\sum_{i=1}^{d}t(k_{i})\,,$
where $k_{i}$ denotes the number of vertices in $V_{i}$. By linearity of $t$,
$\overline{k}=\frac{1}{d}\sum_{i=1}^{d}k_{i}\implies t(\overline{k})=\frac{1}{d}\sum_{i=1}^{d}t(k_{i})=\frac{1}{d}T(G,\mathcal{SC}_{sys}).$
Also assuming that $d$ is minimum,
$T(G,\mathcal{SC}_{sys})=\mathbb{c}_{2}(G)\cdot t(\overline{k})\,.$
It is possible that $T(G,\mathcal{SC}_{sys})$ can be minimized by some $\mathcal{SC}_{sys}$ that is not of minimum size $\mathbb{c}_{2}(G)$. However, the case $d=\mathbb{c}_{2}(G)$ provides an upper bound $T(G,\mathcal{SC}_{sys})=\mathbb{c}_{2}(G)\cdot t(\overline{k})$ on distribution time, which we conjecture is a tight bound for most graph states.
Given a particular $\mathcal{SC}_{sys}$, we can also optimize the ordering. Note that we do not always have to do a complete edge reset, there are some edges that we could potentially keep, minimizing both the number of gates and time steps needed.
Another useful result [11] is as follows.
Theorem IV.2.
For any graph $G$ with $n$ vertices,
$\mathbb{c}_{2}(G)\leq n-1\,.$
If $c_{2}(G)$ is constant with respect to some class of graphs, then the graph states corresponding to that class could be efficiently distributed in a linear number of time steps since $t(\overline{k})\leq n$, so $T(G,\mathcal{SC}_{sys})=O(n)$ (from IV.1). It turns out we can characterize the class of graphs with a constant bounded value of $c_{2}(G)$. Formally, $\{G|\mathbb{c}_{2}(G)\leq d\}$ can be characterized by a finite number of forbidden induced subgraphs (i.e. there is no subgraph in $G$ isomorphic to any graph in some finite set) for all $d\in\mathbb{N}$ [10]. Dense structured graphs often have a bounded $d$, such as cliques ($d=1$) and complete bipartite graphs $(d=3)$. Specifically, the $\mathcal{SC}_{sys}$ with minimum $T(G,\mathcal{SC}_{sys})$ for cliques just corresponds to one $SC$ on all vertices. Similarly, for complete bipartite graphs with vertex partitions $P_{1},P_{2}$ the corresponding optimal subgraph complementation system is given by $\mathcal{SC}_{sys}=\{P_{1}\cup P_{2},P_{1},P_{2}\}$.
[…]
Calculating the optimal number of time steps taken to distribute a graph state and the corresponding $\mathcal{SC}_{sys}$ appears to be a hard problem in general, but we observe some common characteristics of entanglement distribution using subgraph complementations. There is often a trade-off between $t(\overline{k})$ and $\mathbb{c}_{2}(G)$, for some $\mathcal{SC}_{sys}$, such that their product $T(G,\mathcal{SC}_{sys})=O(n)$ for structured or common classes of graphs. Intuitively, if larger subgraphs are complemented at each step, then fewer steps are needed to build an arbitrary structured graph, and vice-versa. We also note that greedy vertex cover algorithms provide good constant factor approximations for $\mathcal{SC}_{sys}$ w.r.t. $\mathbb{c}_{2}(G)$ [10], and hence also $T(G,\mathcal{SC}_{sys})$. Table I lists the asymptotic costs for several relevant classes of graphs.
TABLE I: Costs for distributing various graph states
Class $\mathbf{\mathbb{c}_{2}(G)}$ $\mathbf{t(\overline{k})}$ $\mathbf{T(G,\mathcal{SC}_{sys})}$
Sparse Graphs: $O(n)$ $O(1)$ $O(n)$
Path Graph $n-1$ $O(1)$ $O(n)$
Cycle Graph $n-2$ $O(1)$ $O(n)$
Tree Graph $\leq n-2$ $O(1)$ $O(n)$
Dense Graphs: $O(n)$
Complete Graph ($n$-GHZ state) $1$ $O(n)$ $O(n)$
Repeater Graph [12] $\leq 3$ $O(n)$ $O(n)$
Complete $m$-partite Graph $m+1$ $O(n)$ $O(n)$
IV-C Entanglement of the graph state
An interesting property of $\mathbb{c}_{2}(G)$ is its relation to the Schmidt-rank $r_{G}(X)$, a common entanglement measure of pure states. For a graph state $\left|G\right\rangle$, $r_{G}(X)$ is defined over some bipartition $(X,V(G)\setminus X)$ of the vertices of $G$. The Schmidt-rank quantifies the number of superpositions of a pure state into factorizable states [13] —- a higher Schmidt-rank of a state implies it has higher entanglement. Define the average Schmidt-rank $\mathbb{E}_{r}(G)$ to be the expectation of $r$ for a uniformly chosen random subset of $V$.
Theorem IV.3.
For any graph $G$, $\mathbb{c}_{2}(G)>\mathbb{E}_{r}(G)\,.$
Proof.
This can be inferred from a relation between $\mathbb{c}_{2}(G)$ and the cut-rank of a graph [14], which equals the Schmidt-rank [15], as described below. Define the cut-rank $\rho_{G}(X)$ of a set $X$ of vertices in a graph $G$ as the rank of the $X\times(V(G)\setminus X)$ matrix over the binary field whose $(i,j)$-entry is 1 if vertex $i$ in $X$ is adjacent to vertex $j$ in $V(G)\setminus X$ and 0 otherwise. Define the expected cut-rank $\mathbb{E}\rho(G)$ as the cut-rank for a uniformly chosen random subset $X$ of $V(G)$ [14]. Now, consider the following two results:
$\mathbb{c}_{2}(G)$ $>\mathbb{E}\rho(G)\,,$ (1)
$\rho_{G}(X)$ $=r_{G}(X)\,,$ (2)
where (1) comes from [14] and (2) from [15]. From these, we can directly conclude that $\mathbb{E}\rho(G)=\mathbb{E}_{r}(G)$, and thus,
$\mathbb{c}_{2}(G)>\mathbb{E}_{r}(G)\,.$
Consequently, graph states with a high degree of entanglement will have large values of the subgraph complementation number $\mathbb{c}_{2}(G)$, and therefore may be more expensive to distribute.
V Comparison to prior work
Our approach utilizes $(n-1)$ EPR pairs, $O(\overline{k}n)$ bits of classical communication (CC), and a maximum of $n-1$ qubits at the central node, all of which improve on prior work [2, 3, 8]. Our approach requires one time step to distribute an arbitrary graph state, if we follow early work to implement our cost model [2, 16], which make the assumption that local gates at multiple nodes and the same Clifford operations on a single central node commute, so all the operations can be performed in one time step. We instead assume a gate-based model of computation, which implies $T(G,\mathcal{SC}_{sys})$ number of time steps is required. Other work [8] utilizes an approach with a central network node that takes $O(n)$ time steps if local operations are free and $O(|E|)=O(n^{2})$ time steps if not (accounting for local gates) to build a graph state corresponding to a dense graph. In the presence of memory decoherence, local operations are not free. The fidelity of the state is directly dependent on the number of time steps required. Our approach may also utilize $O(n^{2})$ number of gates and time steps to distribute certain dense graph states but not for most of the common classes since $T(G,\mathcal{SC}_{sys})=O(n)$ even for several classes of dense graphs. This presents a quadratic improvement in the number of time steps, gates required, and potential fidelity of the state in the presence of noise.
VI Incorporating Parallelization
VI-A Log depth GHZ state generation
We first describe a log depth circuit to generate a GHZ state, that is the same up to local rotations as the construction proposed by Cruz et al. [17]. This will be useful as a subroutine in reducing the time to create a star graph in steps 1 and 3 of the subgraph complementation protocol from III-A.
Let $n$ qubits be denoted by $q_{0},q_{1},q_{2}...q_{n-1}$. Begin with a GHZ state (star graph state with center $q_{0}$) of a constant size of $c$ qubits s.t. $c<n/2$ which can be created in constant depth through the application of $c$ $CZ$ gates. We claim that the size of the GHZ state can be doubled in constant depth. This can be done as follows.
  1. 1.
    $\forall q_{i}:1\leq i\leq c\implies CZ(q_{i},q_{c+i})$. The gates can be carried out in parallel, taking one time step.
  2. 2.
    $\forall q_{i}:1\leq i\leq c\implies LC(q_{i})$. Also takes one time step.
  3. 3.
    $\forall q_{i}:1\leq i\leq c\implies CZ(q_{i},q_{c+i})$. This is identical to step 1, thus it also takes one time step.
The above protocol is equivalent to applying $CZ(q_{0},q_{c+i}),\forall q_{i}:1\leq i\leq c$ but only takes three time steps (instead of $c$ time steps). Thus, the size of the GHZ state is doubled from $c$ to $2c$ in constant time. Repeated application of this protocol clearly leads to the creation of an $n$-qubit GHZ state in $O(\log(n))$ depth.
VI-B Application to subgraph complementation
Now, the above protocol can be utilized to reduce the times for steps 1 and 3 of the subgraph complementation protocol from Section III-A. We first observe that in both of those steps, a star graph is being added (modulo 2) on the induced subgraph of the qubits $\forall i:a_{i}$ with $a_{0}$ as the center.
Let there be $n$ qubits in the central node. Naively, it takes $n-1$ $CZ$ gates (and time steps) to add the star graph. We claim that by using $k$ auxiliary qubits $q_{0},q_{1}...q_{k-1}$ in the central node, the star graph can be added in $O(n/k+\log(k))$ time steps. Since steps 2 and 4 of the subgraph complementation protocol both take one time step, this reduces the asymptotic runtime for carrying out an $SC$ operation to the time required to add a star graph (in steps 1 and 3). The protocol is described as follows.
Figure 4: Parallelized subgraph complementation. The solid lines indicate the inner edges whereas the dashed lines indicate the outer edges which were added in step (c) of the parallelized $SC$ operation protocol.
Parallelized SC operation protocol for steps 1 and 3
  1. (a)
    Create a $k$-qubit GHZ state at the central node on the auxiliary $k$ qubits by using the protocol from Section VI-A. This takes $O(\log(k))$ time steps.
  2. (b)
    Add $(k-1)$ edges of the form $(q_{0},a_{i})$ through the doubling protocol described in section VI-A . This can be done on any subset of size $(k-1)$ of qubits $a_{i}$.
  3. (c)
    Repeat step (b) $n/(k-1)$ times to create a star graph with $(k-1)$ inner edges (from the initial $k$-qubit GHZ state) and $n$ outer edges (between $q_{0}$ and $a_{i}$). Fig. 4 illustrates this construction.
  4. (d)
    Remove the $(k-1)$ inner edges, similar to step (a) This also takes $O(\log(k))$ time steps.
Note that step 4 of the original subgraph complementation protocol now becomes $LC(q_{0})$ instead of $LC(a_{0})$.
VI-C Analysis
Let $T(G)$ denote the time required by the parallelized protocol to distribute a graph state $\left|G\right\rangle$. Denote by $t(n)$ the time required for a single subgraph complementation on $n$ vertices, $n\in\mathbb{N}$. Steps 1 and 4 of the above protocol take $O(\log n)$ time. Steps 2 and 3 take $O(n/k)$ time. The other steps in the original subgraph complementation protocol take constant time. Thus, with $(n+k)$ qubits in the central node,
$t(n)=O\left(\frac{n}{k}+\log(k)\right).$
Choosing $k=n/\log(n),$
$t(n)=O(\log(n)).$
Consequently,
$T(G)=\mathbb{c}_{2}(G)\cdot t(\overline{n})=O(n\log(n))\,.$ (3)
We can show that a tighter bound holds with high probability for the distribution of random graphs. The minimum rank of random graph over the field $\mathbb{F}_{2}$ is bounded according to the following theorem [18].
Theorem VI.1.
For an Erdős-Rényi random graph $\mathcal{G}_{n,p}$, and any constant $10<p<$ 1, it holds with high probability that
$\mathbb{c}_{2}(\mathcal{G}_{n,p})=\Theta\left(\frac{n}{\log n}\right).$
Thus, since $t(n)=O(\log n)$ by the prior result, we have with high probability that
$T(\mathcal{G}_{n,p})=\mathbb{c}_{2}(\mathcal{G}_{n,p})\cdot t(\overline{n})=O\left(\frac{n}{\log n}\right)\cdot O(\log n)=O(n)\,.$ (4)
for a random graph $\mathcal{G}_{n,p}$.
VII Graph State Generation in the Presence of Noise
In this section, we present numerical results comparing our graph state distribution algorithm to the algorithm proposed by Avis et al. [8], which is closest to ours in terms of problem setting and resource consumption  [^1] All the code used to carry out simulations, including programs to create the graphs in this sections, are publicly available at https://github.com/Anisen123/graph-state-simulator..
To carry out our simulations, we model noise through depolarizing channels. The action of a depolarizing channel on a state $\rho$ can be expressed by [19]:
$\mathcal{E}(\rho)\xrightarrow{}(1-p)\rho+\frac{p}{3}(X\rho X+Y\rho Y+Z\rho Z)\,.$ (5)
This is equivalent to applying $X,Y$ or $Z$ gates on $\rho$ with equal probability $p/3$ and the identity $I$ with probability $(1-p)$. Each time an operation ($LC$ or $CZ$) is carried out on some qubit $q_{i}$ with state $\rho_{i}$, we model noise by applying 5 on the qubit. Thus, the graph state $\left|G\right\rangle$ is transformed as follows:
$\left|G\right\rangle\xrightarrow{}P_{i}\left|G\right\rangle\,,$
where the Pauli operator $P_{i}\in\{X,Y,Z,I\}$ acts on the state $\rho_{i}$ and $P_{i}$ is selected according to probabilities described above. We can also express $P_{i}\left|G\right\rangle$ as
$P_{i}\left|G\right\rangle=Z^{b}\left|G\right\rangle\,,$
for some bitstring $b=b_{1}b_{2}...b_{n}$, where $Z$ is applied on the $i$’th qubit iff $b_{i}=1$, since the action of any Pauli operator on $q_{i}$ can be mapped to $Z$ gates acting on some set of qubits in the graph state [20]. Thus, we can now express all noise acting on $\left|G\right\rangle$ as a single bitstring of $Z$ operators acting on $\left|G\right\rangle$ since $Z$ gates commute. This makes the noise easier to characterize and store in a list data structure. We estimate the fidelity by running this simulation several times (between 5000 and 10000 times) and finding the fraction of states where $b$ is the all zero bitstring. Note that the probability that $b$ is the all zero bitstring corresponds to the fidelity of the total mixed state. This is because the set of all possible states of the form $Z^{b}\left|G\right\rangle$ corresponding to $2^{n}$ possible linearly independent $Z$-bitstrings are all orthogonal to each other.
The red line in each simulation graph indicates an arbitrarily chosen fidelity of 0.75, which serves as a reference line or a visual aid to compare fidelities between different graphs.
VII-A GHZ State Generation
Fig. 5 compares the fidelity obtained using our protocol with that obtained using the protocol by Avis et al. [8] to generate a 3-qubit GHZ state, as a function of probability $p$ of applying a non-identity Pauli gate at each operation. We note that we implement the local creation of a GHZ state (for [8]) by creating a star graph state first and then applying an $LC$ operation on the central node. Fig. 5 shows that our algorithm distributes small GHZ states with higher fidelity. Fig. 6 compares fidelities varying the number of qubits for a fixed $p=10^{-4}$, and we observe that there is a larger difference in fidelity to distribute graph states with a higher number of qubits.
Figure 5: Comparison of fidelity of a 3-qubit GHZ state distributed according to the subgraph complementation algorithm and the factory node + teleportation algorithm [8] having $0\leq p\leq 0.015$, with the results obtained over 10000 trials. The red line indicates a fidelity of 0.75.
Figure 6: Comparison of fidelity of an $n$-qubit GHZ graph state distributed according to the subgraph complementation algorithm and the factory node + teleportation algorithm [8] varying $6\leq n\leq 30$ with $p=10^{-4}$, with the results obtained over 5000 trials. The red line indicates a fidelity of 0.75.
VII-B Complete bipartite graph state generation
Complete bipartite graphs have applications in entanglement purification [21], silicon-photonic quantum computing [22] and are also used as graph states for all-photonic repeaters [12]. Here, we consider complete bipartite graphs with equal size partitions. Fig. 7(a) and Fig. 7(b) compare the fidelity obtained by implementing our algorithm with the algorithm of Avis et al. [8] for complete bipartite graph states with 6 qubits and 20 qubits respectively. Our algorithm creates complete bipartite graph states with lower fidelity when the number of qubits is small (less than 6 qubits), but with higher fidelity for larger states with 20 qubits. Fig. 8 compares fidelity when the number of qubits is varied between 6 and 30 (with steps of 2 qubits), for error probability $p=10^{-3}$.
(a) 6-qubit complete bipartite graph state
(b) 20-qubit complete bipartite graph state
Figure 7: Comparison of fidelity of (a) 6-qubit complete bipartite graph state and (b) 20-qubit complete bipartite graph state, distributed according to the subgraph complementation algorithm and the factory node + teleportation algorithm [8] having $0\leq p\leq 0.003$, with the results obtained over 10000 trials for each of the graph states. The red line indicates a fidelity of 0.75.
We expect the difference in fidelity to monotonically increase in favor of the subgraph complementation algorithm as the number of qubits increase. We conjecture this difference is a consequence of the $O(n)$ scaling in time and gate operations for our algorithm as compared to $O(n^{2})$ for [8]. For a smaller number of qubits, a linear overhead due to the edge reset operation results in a higher number of operations to carry out a subgraph complementation system as compared to locally creating a graph state and teleporting all the qubits. However, as the number of qubits increases asymptotically, the fidelity to distribute a complete bipartite graph state for our algorithm should increase only linearly as compared to quadratically for [8]. We observe that Fig. 8 exhibits this trend.
Figure 8: Comparison of fidelity of an n-qubit complete bipartite graph state distributed according to the subgraph complementation algorithm and the factory node + teleportation algorithm [8] varying $6\leq n\leq 30$ with $p=10^{-3}$, with the results obtained over 5000 trials. The red line indicates a fidelity of 0.75.
VII-C Parallelized GHZ state generation
We also carried out simulations of distributing a GHZ state with parallelized gate operations as described in section VI. We compare our parallelized SC operation protocol to a parallelized version of the algorithm by Avis et al., in which a complete graph is created locally in the central node by parallelizing $CZ$ gates according to an edge coloring of the complete graph [23]. This algorithm requires $O(n)$ time steps to distribute an $n$-qubit GHZ state, however the parallelized subgraph complementation protocol only requires $O(\log n)$ time steps. Fig. 9(a) compares the fidelity produced by our parallelized SC algorithm with that produced by the parallelized factory node + teleportation algorithm for a 6-qubit GHZ state, as a function of error probability. The parallelized SC algorithm uses 2 auxiliary qubits.
(a) 6-qubit GHZ state (Parallel Implementation)
(b) 20-qubit GHZ state (Parallel Implementation)
Figure 9: Comparison of fidelity of (a) 6-qubit GHZ state and (b) 20-qubit GHZ state, distributed according to the parallelized subgraph complementation algorithm (with (a) 2 auxiliary qubits and (b) 4 auxiliary qubits respectively) and the factory node + teleportation algorithm having $0\leq p\leq 0.015$, with the results obtained over 10000 trials. The red line indicates a fidelity of 0.75.
Our algorithm provides marginally higher fidelity in creating a 6-qubit GHZ state, however we observe in Fig. 9(b) that this difference increases for distributing a larger GHZ state with 20 qubits having error probability of the order of $10^{-4}$.
This trend clearly shows up in Fig. 10, where the number of qubits varies between 2 and 30 to create the corresponding GHZ state.
Figure 10: Comparison of fidelity of an $n$-qubit GHZ graph state distributed according to the parallelized subgraph complementation algorithm and the factory node + teleportation algorithm varying $6\leq n\leq 30$ with $p=10^{-4}$, with the results obtained over 5000 trials. The red line indicates a fidelity of 0.75.
It is interesting to note that on comparing Fig. 6 and Fig. 10, we observe that the fidelity obtained from distributing $n$-qubit GHZ states without parallelization is higher than with parallelization (although both outperform the parallelized factory node + teleportation algorithm). This is expected since we are using $n/\log(n)$ extra qubits for the parallelized protocol. However, we are also gaining an exponential decrease in distribution time ($O(\log n)$ with parallelization as compared to $O(n)$ without it). Therefore, this tradeoff would be favourable in most cases, especially when considering memory decoherence of the qubits which would increase proportionally with distribution time.
VIII Conclusion
We have presented a novel and efficient protocol for distributing an arbitrary graph state across a quantum network by drawing upon the mathematical theory of subgraph complementations, which we show to be strongly linked to the entanglement and structure of the graph state. We show, through simulations, that this protocol generates graph states having high fidelity in the presence of noisy operations. This algorithm also provides a natural framework to create a graph state in a distributed fashion, since subgraph complementations commute and it may be possible to implement several $SC$ operations in parallel (since their order does not matter) and merge the subgraphs at the end.
This could also be extended to a procedure for distributing arbitrary stabilizer states, since graph states are equivalent to stabilizer states up to local operations. We anticipate that the framework of subgraph complementations will be useful for other applications as well, for instance quantum error correction and secret sharing. Further work includes incorporating error correction in estimating runtime and fidelity through noisy simulations of our algorithm, considering the detailed impact of memory decoherence, as well as optimizing our protocol for more domain-specific purposes.
Appendix A Additional simulation results
Figure 11: Comparison of fidelity of a 5-qubit (top) and 20-qubit (bottom) GHZ state distributed according to the subgraph complementation algorithm and the factory node + teleportation algorithm [8] having $0\leq p\leq 0.0015$, with the results obtained over 10000 trials. The red line indicates a fidelity of 0.75.
Figure 12: (Simulation considers noisy gate operations and memory decoherence) Comparison of fidelity of a 3-qubit GHZ state distributed according to the subgraph complementation algorithm and the factory node + teleportation algorithm [8] having $0\leq p\leq 0.0015$, with the results obtained over 10000 trials. The red line indicates a fidelity of 0.75.
Figure 13: (Simulation considers noisy gate operations and memory decoherence) Comparison of fidelity of an $n$-qubit GHZ graph state distributed according to the subgraph complementation algorithm and the factory node + teleportation algorithm [8] varying $6\leq n\leq 30$ with $p=10^{-4}$, with the results obtained over 5000 trials. The red line indicates a fidelity of 0.75.